OpenStudy (anonymous):

What are the solution(s) to the following system of equations? y = x2 – 1 y = x – 1

OpenStudy (anonymous):

0,-1 and -1,0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Set the right sides of the equations equal to each other. It will become a rather simplified quadratic.

OpenStudy (anonymous):

0,-1 and 1,0

OpenStudy (anonymous):

You got it! Good job!

OpenStudy (anonymous):

how about this one

OpenStudy (anonymous):

Which system of equations matches the graph below? graph of a line starting at a closed circle at (1, negative 2) and passing through (2, 0). The second graph is a partial parabola starting at an open circle at (1, 0) and continuing through point (negative 1, 0) y = x2 – 1 if x < 1 y = 2x – 4 if x greater than or equal to –2 y = x2 – 1 if x > –1 y = 2x – 4 if x greater than or equal to –2 y = x2 – 1 if x < 1 y = 2x – 4 if x greater than or equal to 1 y = x2 – 1 if x < 0 y = 2x – 4 if x greater than or equal to –2

OpenStudy (anonymous):

Can you put this in a new post? I might not have time to stick around and do it and if you leave it here, you might not get anyone to come by because they'll look at the first question, see it answered, and leave.